OpenStudy (anonymous):

A car is pulled with force of 10000N. The cars mass is 1267 kg but the car covers 394.6 in 15 seconds. What is expected acceleration of the car from 10000 N force?

OpenStudy (midhun.madhu1987):

Force = mass x acceleration So, Acceleration= Force/Mass i think so....

OpenStudy (cwrw238):

yes that's right - Newtons Second Law of motion

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

What is the actual acceleration of the car from the observed data of x and t?

OpenStudy (cwrw238):

assuming the car starts from rest and the acceleration is constant for the whole distance then 394.6 = 0.5 * a * 15^2 where a = accelaration

OpenStudy (cwrw238):

a = 394.6 / 112.5